OpenStudy (anonymous):

Please Help! If a coin is flipped 10 times, what is the probability that it will show all heads or all tails? 1 2 1 512 1 1,022 1 1,024 1 2,048

OpenStudy (anonymous):

i think one half

OpenStudy (anonymous):

1/2 i believe

OpenStudy (anonymous):

1/512

OpenStudy (anonymous):

well, based on simple probability, we know that it is 1/2 for the first flip. but for both to be the same? 1/4. We keep going like this 3rd flip - 1/8 4th flip - 1/16 5th flip - 1/32 6th - 1/64 7th - 1/128 8th - 1/256 9th - 1/512 10th - 1/1024 so for all the flips to be either heads or tails is a 1/1024 chance

OpenStudy (anonymous):

for 10 heads, the probability=1/2^10=1/1024 for 10 tails, same, 1/1024 so 10 heads or 10 tails->1/1024+1/1024=1/512

OpenStudy (anonymous):

@Curycoo explained it very well

OpenStudy (anonymous):

there is a good vlogbrothers video that i think explains this exact question if you have acess to youtube, too, just search "Vlogbrothers probability" and you'll get a good explanation